An indoor heater operating on the Carnot cycle is warming the house up at a rate of 30 kJ/s to maintain the indoor temperature at 72 ºF. What is the power operating the heater if the outdoor temperature is 30 ºF?


1
Expert's answer
2021-10-31T18:15:43-0400

"T_2 = 30\; °F = 272.039\; K"

Efficiency of heater "= 1- \frac{T_2}{T_1} = 1- (\frac{272.039}{295.372}) = 0.0789953"

The power operating the heater "= \frac{30 \; kJ/s }{ 0.0789953}"

"= 379.77 \; kJ/s \\

= 380 \; kJ/s"